# Duals and time-reversed flows of generalized Ornstein-Uhlenbeck processes

### v3

@@ 1 Introduction
Nowadays, the duality relation (1.1) is widely known and commonly used in applied probability. To illustrate its applicability, assume for the moment being that $(Y_{t}^{y})_{t\geq 0}=(R_{t}^{y})_{t\geq 0}$ is a Cramér-Lundberg risk process absorbed in ruin. This process can be defined by
[…]
@@
see [17] for the earliest reference on this relation, or [2, Chapter III.2] for a textbook treatment. (Actually, apart from [1, Example 3.5], where no proof is given, in most references equation (1.1) is only verified for the case $y\geq 0$ and $x=0$. However, the proof given in [2, Chapter III.2] can be easily adapted to work also for general $x,y\geq 0$). In the present situation, the ruin probability can be written as
@@
where $V$ is a generic random variable whose law is given by the stationary distribution of the dual process $(V_{t})_{t\geq 0}$ which is assumed to exist and to have no atom at $x$, cf. [2, Chapter VII.7].
@@
Thus, in this situation, duality allows to transform the hitting (or ruin) probability of $(R_{t}^{y})_{t\geq 0}$ into the cdf of the stationary distribution of $(V_{t})_{t\geq 0}$, and vice versa. This fact has been used in order to derive eclectic results both in risk theory and queueing theory, see [2, Chapter VII.7] and references therein.
Further examples of pairs of dual Markov processes in the sense of (1.1) can be found e.g. in [1, 10, 21].
Embedding the Cramér-Lundberg process $(R_{t}^{y})_{t\geq 0}$ into an investment market whose dynamics are driven by a Lévy process, say $(W_{t})_{t\geq 0}$, yields a capital process that is often referred to as Paulsen’s risk model [16]. This is a risk process $(R_{t}^{y})_{t\geq 0}$ whose dynamics are described by the stochastic differential equation (SDE)
$\mathop{}\!\mathrm{d}R_{t}^{y}=R_{t-}^{y}\mathop{}\!\mathrm{d}W_{t}+\mathop{}\!\mathrm{d}K_{t},\quad t\geq 0,\quad R_{0}^{y}=y.$ (1.2)
Assuming that $W$ has a jump part of finite variation, in [16] the solution of (1.2) is studied and also an expression for its ruin probability is derived.
In this paper, we treat Siegmund-duality for (time-homogeneous) Markov processes with state space $\mathbb{R}$, defined by (1.1) when $t\geq 0$ and arbitrary $x,y\in\mathbb{R}$ are allowed. We consider solutions of the SDE $\mathop{}\!\mathrm{d}V_{t}^{x}=V_{t-}^{x}\,\mathop{}\!\mathrm{d}U_{t}+\mathop{}\!\mathrm{d}L_{t}$, $V_{0}^{x}=x$, for a general bivariate driving Lévy process $(U_{t},L_{t})_{t\geq 0}$. Assuming that $U$ does not admit jumps of size $-1$ these solutions are known as generalized Ornstein-Uhlenbeck (GOU) processes, see [6, Thm. 2.1], and they can be expressed explicitly. We use this fact to prove in Section 3 that the dual of a GOU process $(V_{t}^{x})_{t\geq 0}$, if existent, is again a GOU process, namely the GOU process $(R_{t}^{y})_{t\geq 0}$ defined by (1.2). We furthermore discuss the relation between hitting probabilities and stationary distributions of the dual pair of GOU processes. As it turns out, the appearing (causal) stationary distribution of the dual GOU process coincides with the non-causal stationary distribution of the original GOU process. When $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ can also be viewed as a $[0,\infty)$-valued Markov process when only $x\geq 0$ is allowed, while the derived dual GOU process $(R_{t}^{y})_{t\geq 0}$ on all of $\mathbb{R}$ is driven by the negative of a subordinator, hence becomes also negative, even when $y\geq 0$. Hence, the Siegmund-dual process $\widehat{R}_{t}^{y}$ of $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$, viewed as an $[0,\infty]$-valued process, must be different from the Siegmund-dual GOU process $(R_{t}^{y})_{t\geq 0}$, viewed as a Markov process on $\mathbb{R}$. In Proposition 3.9 we clarify the relation between $R$ and $\widehat{R}$.
As it is often the case for dual processes, there is an intrinsic relation between time-reversal of a Markov process (more precisely, the stochastic process induced by the reversed stochastic flow) and its dual process, see e.g. [21, 13], and even the duality relation of the Cramér-Lundberg risk process absorbed in ruin and the M/G/1-queue workload process as described above is proved in [2, Chapter III.2] via a time-reversal argument. Motivated by this, we turn our attention to time-reversed GOU processes in Section 4 and prove that the process associated to the reversed stochastic flow is again a GOU process, and - moreover - that it provides a version of the dual GOU process.
2 Preliminaries
General information regarding Lévy processes can be found e.g. in [7, 8, 19], and for stochastic integration we refer to [18]. Consider the stochastic differential equation (SDE)
$\mathop{}\!\mathrm{d}V_{t}^{x}=V_{t-}^{x}\mathop{}\!\mathrm{d}U_{t}+\mathop{}\!\mathrm{d}L_{t},\;t\geq 0,\quad V_{0}^{x}=x,$ (2.1)
where $(U_{t},L_{t})_{t\geq 0}$ is a bivariate Lévy process with characteristic triplet $(\gamma_{U,L},\Sigma_{U,L},\nu_{U,L})$ for some (standard) location parameter $\gamma_{U,L}\in\mathbb{R}^{2}$, a non-negative definite Gaussian covariance matrix $\Sigma_{U,L}=\left(\begin{smallmatrix}\sigma_{U}^{2}&\sigma_{U,L}\\ \sigma_{U,L}&\sigma_{L}^{2}\end{smallmatrix}\right)\in\mathbb{R}^{2\times 2}$, and a Lévy measure $\nu_{U,L}$ on $\mathbb{R}^{2}\setminus\{0\}$. The characteristic function of $(U_{t},L_{t})$ at $(x,y)^{T}\in\mathbb{R}^{2}$ is then given by
$\mathbb{E}\mathrm{e}^{\mathrm{i}(xU_{t}+yL_{t})}=$ $\exp\Big{\{}t\Big{(}\mathrm{i}(x,y)\gamma_{U,L}-(x,y)\Sigma_{U,L}(x,y)^{T}$
$\quad+\int_{\mathbb{R}^{2}\setminus\{0\}}(\mathrm{e}^{\mathrm{i}(x,y)z}-1-\mathrm{i}(x,y)z\mathds{1}_{|z|\leq 1})\,\nu_{U,L}(\mathop{}\!\mathrm{d}z)\Big{)}\Big{\}}.$
We assume that $\nu_{U,L}(\{-1\}\times\mathbb{R})=0$, i.e. $\Delta U_{t}\neq-1$ for all $t\geq 0$. Hereby, for any càdlàg process $Z$ we denote by $Z_{t-}$ the left-hand limit of $Z$ at time $t\in(0,\infty)$, set $Z_{0-}:=Z_{0}$, and write $\Delta Z_{t}=Z_{t}-Z_{t-}$ for its jumps. The marginal jump measures of $U$ and $L$ are denoted by $\nu_{U}$, and $\nu_{L}$, respectively.
Under this assumption it has been shown in [6, Thm. 2.1] that the solution $(V_{t}^{x})_{t\geq 0}$ of (2.1), the generalized Ornstein-Uhlenbeck (GOU) process, is given by
$V_{t}^{x}=\mathcal{E}(U)_{t}\Big{(}x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\Big{)}.$ (2.2)
Hereby, $(\mathcal{E}(U)_{t})_{t\geq 0}$ is the stochastic exponential of $(U_{t})_{t\geq 0}$, i.e. the unique solution of the SDE
which can be expressed explicitly by the Doléans-Dade formula (see [18, Thm. II.37]) as
$\mathcal{E}(U)_{t}=\mathrm{e}^{U_{t}-\sigma_{U}^{2}t/2}\prod_{0<s\leq t}(1+\Delta U_{s})\mathrm{e}^{-\Delta U_{s}},\quad t\geq 0,$ (2.3)
from which it follows that $\mathcal{E}(U)_{t}>0$ for all $t$ if and only if $\Delta U_{t}>-1$ for all $t$. The Lévy process $(\eta_{t})_{t\geq 0}$ in (2.2) is given by
Observe that $\eta_{t}=L_{t}$ if $U$ and $L$ are independent, and that if $L$ is a subordinator and $\Delta U_{t}>-1$ for all $t$, then also $\eta$ is a subordinator, since then $\sigma_{U,L}=0$ and $\Delta\eta_{t}=\Delta L_{t}/(1+\Delta U_{t})$. If $U_{t}=\lambda t$ is chosen deterministically we refer to the resulting GOU process
as Lévy-driven Ornstein-Uhlenbeck process.
Throughout the paper let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ denote the augmented natural filtration of $(U_{t},L_{t})_{t\geq 0}$, and note that by [5, Thm. 3.1] (see also [11]), $(V_{t}^{x})_{t\geq 0}$ is a rich Feller process and hence a Markov process with respect to $\mathbb{F}$.
According to [6, Thm. 2.1], in case of its (almost sure) convergence the law of the exponential functional
$\int_{0}^{\infty}\mathcal{E}(U)_{s-}\mathop{}\!\mathrm{d}L_{s}:=\lim_{t\to\infty}\int_{(0,t]}\mathcal{E}(U)_{s-}\mathop{}\!\mathrm{d}L_{s},$
is the unique causal stationary distribution of $(V_{t})_{t\geq 0}$, while the unique, non-causal stationary distribution of $(V_{t})_{t\geq 0}$ is given by the law of
$-\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}:=-\lim_{t\to\infty}\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathop{}\!\mathrm{d}\eta_{s},$
whenever this converges almost surely. Hereby, a solution of (2.1) with (possibly random) starting point $V_{0}$ is called causal, if $V_{0}$ is independent of $(U,L)$, whereas in case of dependence it is called non-causal. Necessary and sufficient conditions for almost sure convergence of the integrals $\int_{(0,\infty)}\mathcal{E}(U)_{s-}\mathop{}\!\mathrm{d}L_{s}$ and $-\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ have been obtained by Erickson and Maller [9] and are summarised in [6, Thm. 3.5 – Cor. 3.7].
@@ 3 The dual of the generalized Ornstein-Uhlenbeck process
Under certain conditions the class of Lévy-driven OU processes is self-dual in the following sense:
@@
where $(L_{t})_{t\geq 0}$ is a compound Poisson process with non-negative jumps. This process can be interpreted as a Lévy-driven Ornstein-Uhlenbeck process with sticky boundary in zero. According to [1, Example 3.5], the dual of the process $(V_{t}^{x})_{t\geq 0}$ is again a Lévy-driven OU process whose driving process is just $(-L_{t})_{t\geq 0}$ and that is killed when passing 00.
@@ 3.1 Siegmund duality on $\mathbb{R}$
Note that if $\Delta U_{t}>-1$ for all $t$ and $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ has state-space $[0,\infty)$, and hence this GOU process fits directly into the setting originally considered in [20]. For arbitrary choices of $(U,L)$ this is not the case and we shall therefore first consider an extension of Siegmund’s notion of duality to processes on the whole real line as given in [10, Example 1.1], before we come back to the aforementioned special case of processes on the half-line in Section 3.3.
@@
Let $(X_{t}^{x})_{t\geq 0}$ and $(Y_{t}^{y})_{t\geq 0}$ be (universal) time-homogeneous Markov processes on $\mathbb{R}$, where $x,y\in\mathbb{R}$. Then $(Y_{t}^{y})_{t\geq 0}$ is said to be dual (in the sense of Siegmund) to $(X_{t}^{x})_{t\geq 0}$ if $\mathbb{P}(X_{t}\geq y|X_{0}=x)=\mathbb{P}(Y_{t}\leq x|Y_{0}=y)$ for all $0\leq t<\infty$ and $x,y\in\mathbb{R}$, i.e. if for all $x,y\in\mathbb{R}$
@@
Note that it follows immediately from (3.1) that if $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$, then $(-X_{t}^{x})$ is dual to $(-Y_{t}^{y})$. However, duality as defined is not necessarily a symmetric relation as illustrated by the following example.
@@
$\mathbb{P}(X_{t}^{x}\geq y)=\left.\begin{cases}\mathds{1}_{x\geq y},&x\geq 0,\\ \mathbb{P}(N_{t}\leq x-y),&x<0\end{cases}\right\}=\mathbb{P}(Y_{t}^{y}\leq x),$
and hence $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$ in the sense of Definition 3.1. However we note that for $y<0$
@@
    Suppose that $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$. Then a necessary and sufficient condition for $(X_{t}^{x})$ to be dual to $(Y_{t}^{y})$ is that
@@
    Provided that $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$, a sufficient (but not necessary) condition for (3.2) to hold is that $\mathbb{P}(X_{t}^{x}\geq y)$ is continuous as a function in $x$ for each fixed $y\in\mathbb{R}$ and $t>0$, and continuous as a function in $y$ for each fixed $x\in\mathbb{R}$ and $t>0$.
@@
Similar to the original setting of Siegmund-duality described in the introduction, where $[0,\infty]$-valued Markov processes were considered, it is trivial from (3.1) that a necessary condition for an $\mathbb{R}$-valued Markov process $(X_{t}^{x})$ to have a dual process as described in Definition 3.1 is that the function $x\mapsto\mathbb{P}(X_{t}^{x}\geq y)$ is nondecreasing and right-continuous for all $y\in\mathbb{R}$ and $t\geq 0$.
The next theorem shows that the dual of a GOU process $(V_{t}^{x})$ is again a GOU process $(R_{t}^{y})$, and that then $(V_{t}^{x})$ is also dual to $(R_{t}^{y})$. Note that in order to ensure stochastic monotonicity we restrict ourselves to the case $\Delta U_{t}>-1$ for all $t$; see also Remark 3.6 below.
@@
Let $(V_{t}^{x})_{t\geq 0}$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{t},L_{t})_{t\geq 0}$ with $\Delta U_{t}>-1$ for all $t$, as given in (2.2). Then there exists another bivariate Lévy process $(W_{t},K_{t})_{t\geq 0}$ defined by
@@
such that the unique solution $(R_{t}^{y})_{t\geq 0}$ of
is dual to $(V_{t}^{x})_{t\geq 0}$ in the sense of Definition 3.1. Moreover, $(V_{t}^{x})_{t\geq 0}$ is dual to $(R_{t}^{y})_{t\geq 0}$.
@@
Observe first that $\Delta U_{t}>-1$ for all $t$ implies $\mathcal{E}(U)_{t}>0$ for all $t\geq 0$ by (2.3). Thus, using (2.2),
@@
such that in particular $\Delta W_{t}=\frac{-\Delta U_{t}}{1+\Delta U_{t}}>-1$. Moreover, for every $t\geq 0$, by [6, Lemma 3.1],
[…]
for some Lévy process $(Z_{t})_{t\geq 0}$ that can be computed via [6, Eq. (2.2)] as
@@
$-\xi_{t}=\log\mathcal{E}(W)_{t}=W_{t}-\frac{1}{2}\sigma_{W}^{2}t+\sum_{0<s\leq t}\bigg{(}-\Delta W_{s}+\log(1+\Delta W_{s})\bigg{)},\quad t\geq 0,$
using (2.3), we obtain that the SDE solved by $(R_{t}^{y})_{t\geq 0}$ is hence of the form (3.3) with $(W_{t})_{t\geq 0}$ as given in (3.6) and $(K_{t})_{t\geq 0}$ following via [6, Eq. (1.3)] and (3.6) as

### v4

# Duals and inverse flows of generalized Ornstein-Uhlenbeck processes

@@ 2.2 Hitting probabilities
Assume that $(Y_{t}^{y})_{t\geq 0}=(R_{t}^{y})_{t\geq 0}$ is a Cramér-Lundberg risk process absorbed in ruin. This process can be defined by
$R_{t}^{y}:=(y+K_{t})\,\mathds{1}_{\{t<\tau(y)\}},\quad t,y\geq 0,$
@@
see [42] for the earliest reference on this relation, or [1, Chapter III.2] for a textbook treatment. (Actually, apart from [3, Example 3.5], where no proof is given, in most references equation (1.1) is only verified for the case $y\geq 0$ and $x=0$. However, the proof given in [1, Chapter III.2] can be easily adapted for general $x,y\geq 0$). Thus, in the present situation, the ruin probability can be written as
@@
where $V$ is a generic random variable whose law is given by the limiting distribution of the dual process $(V_{t}^{0})_{t\geq 0}$ which is assumed to exist and to have no atom at $y$, cf. [1, Chapter III.2].
@@
Hence, in this situation, duality allows to transform the hitting (or ruin) probability of $(R_{t}^{y})_{t\geq 0}$ into the cdf of the stationary distribution of $(V_{t}^{x})_{t\geq 0}$, and vice versa. This fact has been used to derive various results in both risk theory and queueing theory; see [1, Chapter VII.7] and references therein. Embedding the Cramér-Lundberg process $(R_{t}^{y})_{t\geq 0}$ into an investment market whose dynamics are driven by a Lévy process, say $(W_{t})_{t\geq 0}$, yields a capital process that is often referred to as Paulsen’s risk model [41]. This is a risk process $(R_{t}^{y})_{t\geq 0}$ whose dynamics are described by the SDE
$\mathop{}\!\mathrm{d}R_{t}^{y}=R_{t-}^{y}\mathop{}\!\mathrm{d}W_{t}+\mathop{}\!\mathrm{d}K_{t},\quad t\geq 0,\quad R_{0}^{y}=y,$ (2.13)
which turns out to be of the form (1.2).
Assuming that the driving processes have finite jump activity, in [41] the solution of (2.13) is studied and also an expression for its ruin probability is derived. This result of Paulsen has been generalized to the setting of this paper by Bankovsky and Sly in [4] and then reads as follows.
Proposition 2.7 ([4, Thm. 4]).
Suppose that (2.5) is satisfied and consider the GOU process $(V_{t}^{x})_{t\geq 0}$ solving (1.2). Assume that $\lim_{t\to\infty}\mathcal{E}(U)_{t}^{-1}=0$ almost surely and that the integral $\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ converges almost surely as $t\to\infty$ to a non-degenerate finite random variable $\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ with distribution function $H$. Let $\tau(x)=\inf\{t\geq 0,V_{t}^{x}\leq 0\}$. Then
$\mathbb{P}(\tau(x)<\infty)\,\mathbb{E}[H(-V_{\tau(x)}^{x})|\tau(x)<\infty]=H(-x),\quad x\geq 0,$
where $\mathbb{E}[\cdot|\tau(x)<\infty]$ is interpreted as 0 if $\mathbb{P}(\tau(x)<\infty)=0$.
@@ 3 The dual of the generalized Ornstein-Uhlenbeck process
According to the literature, under certain conditions the class of Lévy-driven OU processes is self-dual in the following sense:
@@
where $(L_{t})_{t\geq 0}$ is a compound Poisson process with non-negative jumps. This process can be interpreted as a Lévy-driven Ornstein-Uhlenbeck process with sticky boundary in zero. According to [3, Example 3.5], the dual of the process $(V_{t}^{x})_{t\geq 0}$ is again a Lévy-driven OU process whose driving process is just $(-L_{t})_{t\geq 0}$ and that is killed when passing 0.
@@ 3.1 Siegmund duality on $\mathbb{R}$
As mentioned in Section 2.1, if $\Delta U_{t}>-1$ for all $t$ and $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ (with $x\geq 0$) has state-space $0,\infty)$, and hence this GOU process fits directly into the setting originally considered in [45]. For arbitrary choices of $(U,L)$ this is not the case and we shall therefore consider an extension of Siegmund’s notion of duality to processes on the whole real line as given in [26, Example 1.1], before we come back to the aforementioned special case of processes on the half-line in Section 3.2.
@@
Let $(X_{t}^{x})_{t\geq 0}$ and $(Y_{t}^{y})_{t\geq 0}$ be (universal) time-homogeneous Markov processes on $\mathbb{R}$, where $x,y\in\mathbb{R}$. Then $(Y_{t}^{y})_{t\geq 0}$ is said to be dual (in the sense of Siegmund) to $(X_{t}^{x})_{t\geq 0}$ if for all $x,y\in\mathbb{R}$
@@
Note that it follows immediately from (3.1) that if $(Y_{t}^{y})_{t\geq 0}$ is dual to $(X_{t}^{x})_{t\geq 0}$, then $(-X_{t}^{x})_{t\geq 0}$ is dual to $(-Y_{t}^{y})_{t\geq 0}$. However, duality as defined is not necessarily a symmetric relation as illustrated by the following example.
@@
$\mathbb{P}(X_{t}^{x}\geq y)=\left.\begin{cases}\mathds{1}_{\{x\geq y\}},&x\geq 0,\\ \mathbb{P}(N_{t}\leq x-y),&x<0\end{cases}\right\}=\mathbb{P}(Y_{t}^{y}\leq x),$
and hence $(Y_{t}^{y})_{t\geq 0}$ is dual to $(X_{t}^{x})_{t\geq 0}$ in the sense of Definition 3.1. However we note that for $y<0$
@@
    Suppose that $(Y_{t}^{y})_{t\geq 0}$ is dual to $(X_{t}^{x})_{t\geq 0}$. Then a necessary and sufficient condition for $(X_{t}^{x})_{t\geq 0}$ to be dual to $(Y_{t}^{y})_{t\geq 0}$ is that
@@
    Provided that $(Y_{t}^{y})_{t\geq 0}$ is dual to $(X_{t}^{x})_{t\geq 0}$, a sufficient (but not necessary) condition for (3.2) to hold is that $\mathbb{P}(X_{t}^{x}\geq y)$ is continuous as a function in $x$ for each fixed $y\in\mathbb{R}$ and $t>0$, and continuous as a function in $y$ for each fixed $x\in\mathbb{R}$ and $t>0$.
@@
Similar to the original setting of Siegmund duality described in the introduction, where $0,\infty]-$valued Markov processes were considered, it is trivial from (3.1) that a necessary condition for an $\mathbb{R}$-valued Markov process $(X_{t}^{x})_{t\geq 0}$ to have a dual process as described in Definition 3.1 is that the function $x\mapsto\mathbb{P}(X_{t}^{x}\geq y)$ is nondecreasing and right-continuous for all $y\in\mathbb{R}$ and $t\geq 0$.
The next theorem characterises when a dual of a GOU process $(V_{t}^{x})_{t\geq 0}$ exists and shows that then the dual is again a GOU process $(R_{t}^{y})_{t\geq 0}$, and that then $(V_{t}^{x})_{t\geq 0}$ is also dual to $(R_{t}^{y})_{t\geq 0}$.
@@
Assume (2.1) and let $(V_{t}^{x})_{t\geq 0}$ be the GOU process given by (2.2). Then a dual process to $(V_{t}^{x})_{t\geq 0}$ in the sense of Definition 3.1 exists if and only if (2.5) is satisfied. In that case, the GOU process $(R_{t}^{y})_{t\geq 0}$ solving the SDE
where $(W_{t},K_{t})_{t\geq 0}$ is another bivariate Lévy process defined by
@@
is dual to $(V_{t}^{x})_{t\geq 0}$. Moreover, $(V_{t}^{x})_{t\geq 0}$ is dual to $(R_{t}^{y})_{t\geq 0}$.
@@
If (2.5) is violated, then $(V_{t}^{x})_{t\geq 0}$ is not stochastically monotone by Lemma 2.5, hence no dual process can exist by Remark 3.4. So suppose from now on that (2.5) is satisfied. Then $\mathcal{E}(U)_{t}>0$ for all $t\geq 0$ by (2.4). Thus, using (2.2),
@@
such that in particular $\Delta W_{t}=\frac{-\Delta U_{t}}{1+\Delta U_{t}}>-1$. Moreover, for every $t\geq 0$, by Lemma 2.2,
$\begin{pmatrix}\mathcal{E}(W)_{t}\\ \int_{(0,t]}\mathcal{E}(W)_{s-}\mathop{}\!\mathrm{d}\eta_{s}\end{pmatrix}\overset{d}{=}\begin{pmatrix}\mathcal{E}(W)_{t}\\ \mathcal{E}(W)_{t}\int_{(0,t]}\mathcal{E}(W)_{s-}^{-1}\mathop{}\!\mathrm{d}Z_{s}\end{pmatrix}$, (3.7)
for some Lévy process $(Z_{t})_{t\geq 0}$ that can be computed via (2.3) as
@@
$\xi_{t}=\log\mathcal{E}(W)_{t}=W_{t}-\frac{1}{2}\sigma_{W}^{2}t+\sum_{0<s\leq t}\bigg(-\Delta W_{s}+\log(1+\Delta W_{s})\bigg),\quad t\geq 0,$
using (2.4), we obtain that the SDE solved by $(R_{t}^{y})_{t\geq 0}$ is hence of the form (3.3) with $(W_{t})_{t\geq 0}$ as given in (3.6) and $(K_{t})_{t\geq 0}$ following via (2.7) and (3.6) as